### v1

@@ 7. Connection to other Laplacians
$\Delta_{S}\mathbf{v}=\Delta_{B}\mathbf{v}+\mathbf{M}\mathbf{v}+\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v},\qquad\Delta_{H}\mathbf{v}=\Delta_{B}\mathbf{v}-\mathbf{M}\mathbf{v},$ (7.1)
where $\mathbf{M}:=\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2}$.
@@
Suppose $\Gamma$ is of class $C^{m+3}$ for some nonnegative integer $m\geq 0$ and let $p\in(1,\infty)$. Then,
@@
    $\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq F_{m,p}(\|\mathbf{f}\|_{m,p,\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{m+1,p;\Gamma}),$
@@
The proofs for ($\textbf{VL}_{H}$), ($\textbf{S}_{H}$), ($\textbf{NS}_{H}$) respectively follow from consecutively applying 4.9, 4.9 and 6.9 with datum $\mathbf{f}\leftarrow\mathbf{f}-\mathbf{M}\mathbf{u}$ as many times as the regularity of $\mathbf{f}$ and $\Gamma$ allows. In turn, the proofs for ($\textbf{S}_{S}$) and ($\textbf{NS}_{S}$) similarly follows from applying 4.9 and 6.9 with data $\mathbf{f}\leftarrow\mathbf{f}+\mathbf{M}\mathbf{u}+\nabla_{\Gamma}g$.
The proof for ($\textbf{VL}_{S}$) is more delicate due to the presence of the $\int_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{u}\,\operatorname{div}_{\Gamma}\mathbf{v}$ term, which is of the same order as the $\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}$ term. For the sake of clarity we express the previously omited variational formulation for ($\textbf{VL}_{S}$); it reads
@@
By 4.3 (Helmholtz decomposition) we can decompose the solution $\mathbf{u}$ of (7.3) as $\mathbf{u}=\mathbf{u}_{\sigma}+\nabla_{\Gamma}\phi$, where $\mathbf{u}_{\sigma}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ and $\phi\in\mathrm{W}^{2,p}_{\#}(\Gamma)$. Similarly any test function $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$ can be written as $\mathbf{v}=\mathbf{v}_{\sigma}+\nabla_{\Gamma}\psi$ where $\mathbf{v}_{\sigma}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)$ and $\psi\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$. Now, choosing $\mathbf{v}=\mathbf{v}_{\sigma}$ in (7.3) yields
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}_{\sigma}:\nabla_{\Gamma}\mathbf{v}_{\sigma}+\int_{\Gamma}\nabla_{\Gamma}\nabla_{\Gamma}\phi:\nabla_{\Gamma}\mathbf{v}_{\sigma}=\int_{\Gamma}(\mathbf{f}+\mathbf{M}\mathbf{u})\cdot\mathbf{v}_{\sigma},\qquad\forall\mathbf{v}_{\sigma}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$ (7.4)
Since $\operatorname{div}_{\Gamma}\mathbf{v}_{\sigma}=0$, we can integrate by parts and use 2.1 to rewrite the second term in the left-hand side as $\int_{\Gamma}\nabla_{\Gamma}\nabla_{\Gamma}\phi:\nabla_{\Gamma}\mathbf{v}_{\sigma}=-\int_{\Gamma}\mathbf{M}\nabla_{\Gamma}\phi\cdot\mathbf{v}_{\sigma}$, thus arriving at
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}_{\sigma}:\nabla_{\Gamma}\mathbf{v}_{\sigma}=\int_{\Gamma}(\mathbf{f}+\mathbf{M}\mathbf{u}-\mathbf{M}\nabla_{\Gamma}\phi)\cdot\mathbf{v}_{\sigma}=\int_{\Gamma}(\mathbf{f}+\mathbf{M}\mathbf{u}_{\sigma})\cdot\mathbf{v}_{\sigma},\qquad\forall\mathbf{v}_{\sigma}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$ (7.5)
Consequently, by 4.9 ($\mathrm{L}^{p}$-based regularity for tangent Stokes), $\mathbf{u}_{\sigma}\in\mathbf{W}^{2,p}_{t}(\Gamma)$ and
@@
Again, by means of 2.1 we can rewrite the first two terms in the left-hand side as
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}_{\sigma}:\nabla_{\Gamma}\nabla_{\Gamma}\psi$ $=-\int_{\Gamma}\mathbf{M}\mathbf{u}_{\sigma}\cdot\nabla_{\Gamma}\psi,$ (7.7)
$\int_{\Gamma}\nabla_{\Gamma}\nabla_{\Gamma}\phi:\nabla_{\Gamma}\nabla_{\Gamma}\psi$ $=\int_{\Gamma}\Delta_{\Gamma}\phi\,\Delta_{\Gamma}\psi-\mathbf{M}\nabla_{\Gamma}\phi\cdot\nabla_{\Gamma}\psi.$ (7.8)
@@
and therefore by 3.22 ($\mathrm{L}^{p}$-based regularity of the biharmonic problem) we deduce that $\phi\in\mathrm{W}^{3,p}_{\#}(\Gamma)$ and
@@ Appendix A Variational problems in reflexive Banach spaces
The relation between the injectivity and surjectivity of the operators $T$ and $T^{\mathrm{t}}$ is given by the following result [12, Section 2.7].
Theorem A.1 (characterization of injectivity and surjectivity).
Let $X$ and $Y$ be two Banach spaces with $Y$ reflexive, and consider $T\in\mathscr{L}(X,Y^{\prime})$. Then, the following set of equivalences hold true
i) $T$ is surjective $\Longleftrightarrow$ $T^{\mathrm{t}}$ is injective and has closed range $\Longleftrightarrow$ there is $\alpha>0$ such that $\sup_{\begin{subarray}{c}x\in X\\ x\neq 0\end{subarray}}\frac{\langle T(x),y\rangle}{\|x\|}\geq\alpha\|y\|,\quad\forall y\in Y$
ii) $T$ is injective and has closed range $\Longleftrightarrow$ $T^{\mathrm{t}}$ is surjective $\Longleftrightarrow$ there is $\widetilde{\alpha}>0$ such that $\sup_{\begin{subarray}{c}y\in Y\\ y\neq 0\end{subarray}}\frac{\langle T(x),y\rangle}{\|y\|}\geq\widetilde{\alpha}\|x\|,\quad\forall x\in X$ .
A direct consequence of A.1 is the celebrated Banach–Nečas–Babuška theorem.
Theorem A.2 (Banach–Nečas–Babuška).
Let $X$ and $Y$ be two Banach spaces with $Y$ reflexive, and consider $T\in\mathscr{L}(X,Y^{\prime})$. Then, $T$ is invertible (and hence $T^{-1}\in\mathscr{L}(Y^{\prime},X)$ by the Bounded Inverse Theorem) if and only if one of the following pair of conditions hold
  1. i)
    If $x\in X$ is such that $\langle T(x),y\rangle=0$ for each $y\in Y$, then $x=0$.
  2. ii)
    There is $\alpha>0$ such that
    $\sup_{\begin{subarray}{c}x\in X\\ x\neq 0\end{subarray}}\frac{\langle T(x),y\rangle}{\|x\|}\geq\alpha\|y\|,\quad\forall y\in Y.$
  1. i)’
    If $y\in Y$ is such that $\langle T(x),y\rangle=0$ for each $x\in X$, then $y=0$.
  2. ii)’
    There is $\widetilde{\alpha}>0$ such that
    $\sup_{\begin{subarray}{c}y\in Y\\ y\neq 0\end{subarray}}\frac{\langle T(x),y\rangle}{\|y\|}\geq\widetilde{\alpha}\|x\|,\quad\forall x\in X.$
Moreover, if $T$ is invertible, then the “optimal” constants $\alpha$ and $\widetilde{\alpha}$ in coincide, that is
$\alpha_{\text{opt}}=\inf_{\begin{subarray}{c}y\in Y\\ y\neq 0\end{subarray}}\sup_{\begin{subarray}{c}x\in X\\ x\neq 0\end{subarray}}\frac{\langle T(x),y\rangle}{\|x\|\|y\|}=\inf_{\begin{subarray}{c}x\in X\\ x\neq 0\end{subarray}}\sup_{\begin{subarray}{c}y\in Y\\ y\neq 0\end{subarray}}\frac{\langle T(x),y\rangle}{\|x\|\|y\|},$
and it also holds that
$\|T^{-1}\|_{\mathscr{L}(Y^{\prime},X)}=\|(T^{\mathrm{t}})^{-1}\|_{\mathscr{L}(X^{\prime},Y)}=\frac{1}{\alpha_{\text{opt}}}.$
The development of the functional analytical framework presented in A.1 and A.2 is the consequence of decades of research built upon the seminal work by Nečas [56, 57], and later enriched by Hayden [38], Babuška [4], Babuška and Aziz [5] and Simader [69]; we refer to [65] for a thorough discussion on its historical development. An exhaustive and self-contained derivation of A.2 can be found in the classical reference [27, Chapter 2] (see also [31, Capítulo 5]).
In several applications the underlying operator $T$ is defined in terms of product spaces and with a “skew-triangular” structure. In this case, A.2 can be equivalently reformulated as the so-called generalized Babuška–Brezzi theorem.
@@
    There is $\alpha>0$ such that
    $\sup_{\begin{subarray}{c}x_{2}\in K_{2}\\ x_{2}\neq 0\end{subarray}}\frac{\langle A(x_{2}),x_{1}\rangle}{\|x_{2}\|}\geq\alpha\|x_{1}\|,\quad\forall x_{1}\in K_{1}.$
@@
    There is $\alpha>0$ such that
    $\sup_{\begin{subarray}{c}x_{1}\in K_{1}\\ x_{1}\neq 0\end{subarray}}\frac{\langle A(x_{2}),x_{1}\rangle}{\|x_{1}\|}\geq\alpha\|x_{2}\|,\quad\forall x_{2}\in K_{2}.$
@@
$\|x_{2}\|$ $\leq\frac{1}{\alpha}\|F\|+\frac{1}{\beta_{2}}\left(1+\frac{\|A\|}{\alpha}\right)\|G\|$
$\|m_{1}\|$ $\leq\frac{1}{\beta_{1}}\left(1+\frac{\|A\|}{\alpha}\right)\|F\|+\frac{\|A\|}{\beta_{1}\beta_{2}}\left(1+\frac{\|A\|}{\alpha}\right)\|G\|.$
A.3 in the Hilbert and symmetric ($B_{1}=B_{2}$) case was proved by Brezzi [13], and later extended to the nonsymmetric case by Nicolaides [58]. A.3 as stated was proved much later by Bernardi, Canuto and Maday [8].
Appendix B Regularity estimates in flat domains
This section contains results on higher Sobolev regularity of elliptic PDEs in flat domains, which are essential for our arguments in Section 3. We were unable to pinpoint a precise reference for these results, but their proof easily follows by an induction argument starting from [35, Lemmas 9.16 and 9.17].
Proposition B.1 (improved differentiability).
Let $\Omega$ be an open connected subset of $\mathrm{R}^{d}$. Let $1<p<\infty$ and an integer $m\geq 0$. Let $u\in\mathrm{W}^{2,p}(\Omega)$ compactly supported on $\Omega$. Moreover, assume that $u$ satisfies
$Lu:=\mathbf{A}:\nabla^{2}u+\mathbf{b}\cdot\nabla u+c\,u=f,\qquad\text{a.e.~on $\Omega$},$ (B.1)
where $\mathbf{A}\in\mathbb{C}^{m}(\overline{\Omega})$, $\mathbf{b}\in\mathbf{W}^{m,\infty}(\Omega)$, $c\in\mathrm{W}^{m,\infty}(\Omega)$, $c\leq 0$, $f\in\mathrm{W}^{m,p}(\Omega)$. Moreover, assume that $\mathbf{A}$ is uniformly elliptic on $\Omega$, i.e. there exists $\Lambda>0$ such that
$\boldsymbol{\xi}\cdot\mathbf{A}(\mathbf{x})\boldsymbol{\xi}\geq\Lambda|\boldsymbol{\xi}|^{2},\qquad\forall\xi\in\mathrm{R}^{d},\quad\forall\mathbf{x}\in\Omega.$
Then $u\in\mathrm{W}^{m+2,p}(\Omega)$ and there exists a positive constant $C$, depending only on $d$, $p$, $\Lambda$, $\mathbf{A}$, $\mathbf{b}$, $c$, $\Omega$ such that
$\|u\|_{m+2,p;\Omega}\leq C\|f\|_{m,p;\Omega}.$
@@ Appendix B Regularity estimates in flat domains
The case $m=0$ follows from [35, Lemma 9.17]. We now proceed by induction. Assume the Lemma holds true for a fixed $m\geq 0$. Consider the setting given by $m$ replaced by $m+1$. Then, by the induction hypothesis $u\in\mathrm{W}^{m+2,p}(\Omega)$ and
$\|u\|_{m+2,p;\Omega}\leq C\|f\|_{m,p;\Omega}.$
For each $i=1,\dots,d$, we differentiate (B.1) to deduce that $w_{i}:=\partial_{i}u\in\mathrm{W}^{m+1,p}(\Omega)$ (which is compactly supported on $\Omega$) satisfies
$\mathbf{A}:\nabla^{2}w_{i}+\mathbf{b}\cdot\nabla w_{i}+c\,w_{i}=g_{i},$ (B.2)
in the sense of $\mathrm{W}^{-1,p}(\Omega)$; where $g_{i}:=\partial_{i}f-(\partial_{i}\mathbf{A}):\nabla^{2}u+(\partial_{i}\mathbf{b})\cdot\nabla u+(\partial_{i}c)\,u\in\mathrm{W}^{m,p}(\Omega)$. By the existence and uniqueness result [35, Theorem 9.15] it follows that $w_{i}\in\mathrm{W}^{2,p}(\Omega)$. Hence, by the induction hypothesis applied to problem (B.2) we deduce that $w_{i}\in\mathrm{W}^{m,2}(\Omega)$ and
$\|\partial_{i}u\|_{m+2,p;\Omega}=\|w_{i}\|_{m+2,p;\Omega}\lesssim\|g_{i}\|_{m,p;\Omega}\lesssim\left(\|f\|_{m+1,p;\Omega}+\|u\|_{m+2,p;\Omega}\right)\lesssim\|f\|_{m+1,p;\Omega}.$
Consequently, $u\in\mathrm{W}^{m+3,p}(\Omega)$ and
$\|u\|_{m+3,p;\Omega}\lesssim\|u\|_{0,p;\Omega}+\sum_{i=1}^{d}\|\partial_{i}u\|_{m+2,p;\Omega}\lesssim\|f\|_{m+1,p;\Omega},$
which finishes the proof. ∎
Proposition B.2 (improved integrability).
In addition to the assumptions of B.1 assume $f\in\mathrm{W}^{m,q}$ for some $q\in(p,\infty)$. Then, the solution $u\in\mathrm{W}^{m+2,p}(\Omega)$ (guaranteed by B.1) lies in $\mathrm{W}^{m+2,q}(\Omega)$ and there exists a positive constant $C$, depending only on $d$, $p$, $q$, $\Lambda$, $\mathbf{A}$, $\mathbf{b}$, $c$, $\Omega$ such that
$\|u\|_{m+2,q;\Omega}\leq C\|f\|_{m,q;\Omega}.$

### v2

@@ 6. Connection to other Laplacians
$\Delta_{S}\mathbf{v}=\Delta_{B}\mathbf{v}+\mathbf{M}\mathbf{v}+\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v},\qquad\Delta_{H}\mathbf{v}=\Delta_{B}\mathbf{v}-\mathbf{M}\mathbf{v},$
where $\mathbf{M}:=\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2}$ is the Ricci curvature tensor.
@@
Suppose $\Gamma$ is of class $C^{m+2,1}$ for some nonnegative integer $m\geq 0$ and let $p\in(1,\infty)$. Then,
@@
    $\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq F_{m,p}(\|\mathbf{f}\|_{m,p;\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{m+1,p;\Gamma}),$
@@
The proofs for ($\textbf{VL}_{H}$), ($\textbf{S}_{H}$), ($\textbf{NS}_{H}$) respectively follow from successive application of 3.7, 4.8 and 5.9 with datum $\mathbf{f}\leftarrow\mathbf{f}-\mathbf{M}\mathbf{u}$ as many times as the regularity of $\mathbf{f}$ and $\Gamma$ permits. Moreover, the proofs for ($\textbf{S}_{S}$) and ($\textbf{NS}_{S}$) similarly follows from applying 4.8 and 5.9 with data $\mathbf{f}\leftarrow\mathbf{f}+\mathbf{M}\mathbf{u}+\nabla_{\Gamma}g$.
The proof for ($\textbf{VL}_{S}$) is more delicate due to the presence of $\int_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{u}\,\operatorname{div}_{\Gamma}\mathbf{v}$, which is of the same order as $\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}$. For the sake of clarity we express the previously omitted variational formulation for ($\textbf{VL}_{S}$); it reads
@@
By 4.3 (Helmholtz–Weyl decomposition on $\Gamma$) we can decompose the solution $\mathbf{u}$ of (6.2) as $\mathbf{u}=\mathbf{u}_{\sigma}+\nabla_{\Gamma}\phi$, where $\mathbf{u}_{\sigma}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ and $\phi\in\mathrm{W}^{2,p}_{\#}(\Gamma)$. Similarly any test function $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$ can be written as $\mathbf{v}=\mathbf{v}_{\sigma}+\nabla_{\Gamma}\psi$ where $\mathbf{v}_{\sigma}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)$ and $\psi\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$. Now, choosing $\mathbf{v}=\mathbf{v}_{\sigma}$ in (6.2) yields
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}_{\sigma}:\nabla_{\Gamma}\mathbf{v}_{\sigma}+\int_{\Gamma}\nabla_{\Gamma}\nabla_{\Gamma}\phi:\nabla_{\Gamma}\mathbf{v}_{\sigma}=\int_{\Gamma}(\mathbf{f}+\mathbf{M}\mathbf{u})\cdot\mathbf{v}_{\sigma},\qquad\forall\mathbf{v}_{\sigma}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$
Since $\operatorname{div}_{\Gamma}\mathbf{v}_{\sigma}=0$, we can use 2.13 (weak commutator identity) to rewrite the second term in the left-hand side as $\int_{\Gamma}\nabla_{\Gamma}\nabla_{\Gamma}\phi:\nabla_{\Gamma}\mathbf{v}_{\sigma}=-\int_{\Gamma}\mathbf{M}\nabla_{\Gamma}\phi\cdot\mathbf{v}_{\sigma}$, thus arriving at
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}_{\sigma}:\nabla_{\Gamma}\mathbf{v}_{\sigma}=\int_{\Gamma}(\mathbf{f}+\mathbf{M}\mathbf{u}-\mathbf{M}\nabla_{\Gamma}\phi)\cdot\mathbf{v}_{\sigma}=\int_{\Gamma}(\mathbf{f}+\mathbf{M}\mathbf{u}_{\sigma})\cdot\mathbf{v}_{\sigma},\qquad\forall\mathbf{v}_{\sigma}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$
Consequently, by 4.8 ($\mathrm{L}^{p}$-based regularity for tangent Stokes), $\mathbf{u}_{\sigma}\in\mathbf{W}^{2,p}_{t}(\Gamma)$ and
@@
Again, by means of 2.13 we can rewrite the first two terms in the left-hand side as
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}_{\sigma}:\nabla_{\Gamma}\nabla_{\Gamma}\psi$ $=-\int_{\Gamma}\mathbf{M}\mathbf{u}_{\sigma}\cdot\nabla_{\Gamma}\psi,$
$\int_{\Gamma}\nabla_{\Gamma}\nabla_{\Gamma}\phi:\nabla_{\Gamma}\nabla_{\Gamma}\psi$ $=\int_{\Gamma}\Delta_{\Gamma}\phi\,\Delta_{\Gamma}\psi-\mathbf{M}\nabla_{\Gamma}\phi\cdot\nabla_{\Gamma}\psi.$
@@
whence by the $\mathrm{L}^{p}$-based regularity of the biharmonic problem [5, §4] we deduce that $\phi\in\mathrm{W}^{3,p}_{\#}(\Gamma)$ and
@@
Acknowledgements
The authors are grateful to professor Giuseppe Savaré (Bocconi University), whose sharp observations on a previous version of this manuscript motivated us to streamline the overall arguments in the paper and improve its results.
@@ Appendix A Variational problems in reflexive Banach spaces
The relation between the injectivity and surjectivity of the operators $T$ and $T^{\mathrm{t}}$ is summarized in [5, Theorem A.1] based on [12, Section 2.7]. Its direct consequence is the celebrated Banach–Nečas–Babuška theorem given in [5, Theorem A.2].
In several applications the underlying operator $T$ is defined in terms of product spaces and with a “skew-triangular” structure. In this case, [5, Theorem A.2] can be equivalently reformulated as the so-called generalized Babuška–Brezzi theorem.
@@
    There is $\alpha>0$ such that sup_x_2 K_2x_2 0 A(x2), x1x2 αx_1,  x_1 K_1.
@@
    There is $\alpha>0$ such that sup_x_1 K_1x_1 0 A(x2), $x1$ αx_2,  x_2 K_2.
@@
$\|x_{2}\|\leq\frac{1}{\alpha}\|F\|+\frac{1}{\beta_{2}}\left(1+\frac{\|A\|}{\alpha}\right)\|G\|\qquad\|m_{1}\|\leq\frac{1}{\beta_{1}}\left(1+\frac{\|A\|}{\alpha}\right)\|F\|+\frac{\|A\|}{\beta_{1}\beta_{2}}\left(1+\frac{\|A\|}{\alpha}\right)\|G\|.$
A.1 in the Hilbert and symmetric ($B_{1}=B_{2}$) case was proved by Brezzi [13], and later extended to the nonsymmetric case by Nicolaides [50]. A.1 as stated was proved much later by Bernardi, Canuto and Maday [7].
Appendix B Calculus identities on manifolds
In this section we present some results on tangential calculus that are instrumental in Section 3. Within this section we adopt the Einstein summation convention.
The following result is used in the proof of 3.6 (well-posedness of the Bochner Laplacian in $\mathbf{W}^{1,p}_{t}(\Gamma)$) to relate, in their weak forms, the Bochner Laplacian operator and the componentwise application of the Laplace–Beltrami operator.
Proposition B.1.
Suppose $\Gamma$ is of class $C^{2}$. Then,
$\nabla_{\Gamma}(\mathbf{P}\mathbf{v})=-(\mathbf{v}\cdot\boldsymbol{\nu})\mathbf{B}+\nabla_{\Gamma}\mathbf{v},\qquad\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}(\mathbf{P}\mathbf{v})=\nabla_{M}\mathbf{u}:\nabla_{M}\mathbf{v}-(\mathbf{v}\cdot\boldsymbol{\nu})\nabla_{M}\mathbf{u}:\mathbf{B}+(\boldsymbol{\nu}\mathbf{u}^{T}\mathbf{B}):\nabla_{M}\mathbf{v},$
for each $\mathbf{u}\in\mathbf{C}^{1}_{t}(\Gamma)$ and $\mathbf{v}\in\mathbf{C}^{1}(\Gamma)$ (not necessarily tangential).
@@ Appendix B Calculus identities on manifolds
We first write,
$\nabla_{\Gamma}(\mathbf{P}\mathbf{v})]_{ij}=[\mathbf{P}\nabla_{M}(\mathbf{P}\mathbf{v})]_{ij}=P_{ik}[\nabla_{M}(\mathbf{P}\mathbf{v})]_{kj}=P_{ik}\underline{D}_{j}(\mathbf{P}\mathbf{v})_{k}=P_{ik}\underline{D}_{j}(P_{kl}v_{l})=P_{ik}\underline{D}_{j}P_{kl}v_{l}+P_{ik}P_{kl}\underline{D}_{j}v_{l}.$
By recalling identity $B_{qr}u_{r}=-\underline{D}_{q}u_{r}\nu_{r}$ (cf. (2.6)), the facts that $\mathbf{B}\mathbf{P}=\mathbf{B}$ and $\mathbf{P}\boldsymbol{\nu}=\mathbf{0}$, and noticing that $\underline{D}_{k}P_{jl}=-B_{kj}\nu_{l}-B_{kl}\nu_{j}$ because $\mathbf{P}_{jl}=\delta_{jl}-\nu_{j}\nu_{l}$ and $\mathbf{B}=\nabla_{M}\boldsymbol{\nu}$, we arrive at
$\nabla_{\Gamma}(\mathbf{P}\mathbf{v})]_{ij}=-P_{ik}B_{kj}\nu_{l}v_{l}-P_{ik}B_{lj}\nu_{k}v_{l}+P_{il}\underline{D}_{j}v_{l}=-(\mathbf{v}\cdot\boldsymbol{\nu})B_{ij}+(\mathbf{P}\nabla_{M}\mathbf{v})_{ij}.$
Hence, it is clear that
$\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}(\mathbf{P}\mathbf{v})=-(\mathbf{v}\cdot\boldsymbol{\nu})\nabla_{M}\mathbf{u}:\mathbf{B}+(\mathbf{P}\nabla_{M}\mathbf{u}):\nabla_{M}\mathbf{v}=-(\mathbf{v}\cdot\boldsymbol{\nu})\nabla_{M}\mathbf{u}:\mathbf{B}+\nabla_{M}\mathbf{u}:\nabla_{M}\mathbf{v}-(\mathbf{N}\nabla_{M}\mathbf{u}):\nabla_{M}\mathbf{v},$
where $\mathbf{N}:=\boldsymbol{\nu}\otimes\boldsymbol{\nu}$. We conclude the proof by noticing that, again by (2.6), it holds that
$(\mathbf{N}\nabla_{M}\mathbf{u}):\nabla_{M}\mathbf{v}=\operatorname*{tr}(\nabla_{M}^{T}\mathbf{u}\,\boldsymbol{\nu}\boldsymbol{\nu}^{T}\nabla_{M}\mathbf{v})=-\operatorname*{tr}(\mathbf{B}\mathbf{u}\,\boldsymbol{\nu}^{T}\nabla_{M}\mathbf{v})=-(\boldsymbol{\nu}\mathbf{u}^{T}\mathbf{B}):\nabla_{M}\mathbf{v}.$
This concludes the proof. ∎
The following result shows that the covariant Hessian $\nabla_{\Gamma}\nabla_{\Gamma}v$ of a scalar function $v\in C^{2}(\Gamma)$ is symmetric, provided $\Gamma$ is of class $C^{2}$.
Proposition B.2 (symmetry of the covariant Hessian).
Suppose $\Gamma$ is of class $C^{2}$. Let the superindex e denote an operator that extends functions defined on $\Gamma$ to a tubular neighborhood $\Omega_{\delta}\subseteq\mathrm{R}^{d+1}$ of $\Gamma$ with the following property: if $v\in C^{2}(\Gamma)$ then $v^{e}:\Omega_{\delta}\rightarrow\mathrm{R}$ is also of class $C^{2}$. Then, identity (2.2) holds true, that is:
[…]